OpenStudy (anonymous):

help with math problem give possible formula for the given graph(will be drawn)

OpenStudy (anonymous):



OpenStudy (mathmath333):

first find the slope (m)

OpenStudy (anonymous):

-9/2

OpenStudy (anonymous):

9/2

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

so formula for the graph i beleive would be y=9/2x+3

OpenStudy (mathmath333):

this one y-y1=m(x-x1)

OpenStudy (anonymous):

wait +c? whats the +c?

OpenStudy (anonymous):

oh kk

OpenStudy (mathmath333):

no sry +c not required

OpenStudy (mathmath333):

in place of (x1,y1) place(0,3) or (2,12)

OpenStudy (anonymous):

hm it tells me that the formula i got is inorrect.

OpenStudy (mathmath333):

in this equation y-y1=m(x-x1)

OpenStudy (anonymous):

thats the graph, its not straight so thats not the correct formula

OpenStudy (mathmath333):

have u put x1 and y1 and m value in y-y1=m(x-x1)?

OpenStudy (anonymous):

yeah